# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

### v2

@@ C.2 Reward Design Details
Table 7: The table for stability analysis. Each method is trained for 5 runs on LibriSpeech and evaluated on the test-clean split. Reborn performs steadily well while being fully converged in both of the stages. The ablations show the effectiveness of applying boundary merging and parameter initialization from the previous iteration for the stage 2 GAN training. The abbreviation "adj." indicates adjacent pooling, which is the 2nd stage of pooling in wav2vec-U (see Appendix D). The "Freq." in the last column is the number of boundaries per second.
C.3 Stability Analysis
Next, we focus on the ablations of Reborn stage 2 GAN training in Table 7. When the boundary merging is not applied, the GAN training on the original Reborn-segmented features is still better than the initial stage. However, the average performance and the standard deviation are larger. The observation indicates that the training is slightly unstable, but does not diverge too much since it is still fully converged. On the other hand, when we train the next-iteration GAN with completely random initialized parameters, the performance degrades even if we train the model for more steps. These two ablations directly demonstrate the effectiveness of adopting boundary merging and parameter initialization from the previous iteration, while the GAN training can still work on the original boundary or with randomly initialized parameters.
C.4 Iteratively Polished Phoneme Predictors
@@
C.5 Generalizability across Different Speech Foundation Models
@@
$\mathcal{L}_{GAN}=$ $\min_{\mathcal{G}}\max_{\mathcal{D}}\underset{Z\sim\mathcal{Z}}{\mathbb{E}}[\log(\mathcal{D}(Z))]+\underset{S\sim\mathcal{S}}{\mathbb{E}}[\log(1-\mathcal{D}(\mathcal{G}(S)))]$ (5)
, where $Z$ is a tokenized phoneme sequence sampled from the phonemicized text corpora $\mathcal{Z}$ and $S$ is the preprocessed speech feature sampled from the preprocessed speech corpora $\mathcal{S}$. Besides the min-max objective in the adversarial training, wav2vec-U [4] also adopts the three regularized objectives for better convergence and performance. First, the gradient penalty $\mathcal{L}_{gp}$ [20] is applied on the discriminator $\mathcal{D}$ to stabilize training, where the input of $\mathcal{D}$ is the linear combination of a generated sample $G$ and a real sample $Z$.
$\mathcal{L}_{gp}=\underset{G,Z,\alpha\sim U(0,1)}{\mathbb{E}}\left[\left(\left\|\nabla\mathcal{D}(\alpha G+(1-\alpha)Z)\right\|_{2}-1\right)^{2}\right]$ (6)
Next, the smoothness penalty $\mathcal{L}_{sp}$ is computed across consecutive segments of a generated logit sequence, promoting the generation of closely aligned neighboring outputs by the model.
$\mathcal{L}_{sp}=\sum_{(g_{t},g_{t+1})\in G}\left\|g_{t+1}-g_{t}\right\|^{2}$ (7)
Furthermore, the phoneme diversity loss $\mathcal{L}_{pd}$ is applied to encourage higher vocabulary usage of the generated sequences. As shown in Eq. 8, the entropy of the averaged softmax distribution of the generator outputs ($H_{\mathcal{G}}(\mathcal{G}(S))$) is maximized over the phoneme vocabulary across a batch B.
$\mathcal{L}_{pd}=\frac{1}{|B|}\sum_{S\in B}-H_{\mathcal{G}}(\mathcal{G}(S))$ (8)
@@
$\mathcal{L}=\mathcal{L}_{GAN}-\lambda\mathcal{L}_{gp}+\gamma\mathcal{L}_{sp}+\eta\mathcal{L}_{pd}$ (9)
@@
Appendix E Implementation Details
E.1 Data Preparation and Phoneme LM Formulation
@@
Our data preparation process mainly follows wav2vec-U [^3] https://github.com/facebookresearch/fairseq/tree/main/examples/wav2vec/unsupervised for fair comparisons. Before the feature extraction stage, we attempt to remove most of the silence through an unsupervised voice activity detection (VAD) method [55]. [^4] https://github.com/zhenghuatan/rVAD However, the method is not completely perfect. To stabilize the GAN training, a special <silence> token is introduced to the text data. The <silence> token is inserted in the front and the end of each sentence, and also between words by a probability of 0.25 according to wav2vec-U. Then, we phonemicize all the words in the text data along with the previously inserted <silence> tokens for GAN training. The two off-the-shelf phonemizers are used to perform phonemicization. For LibriSpeech, we use the G2P phonemeizer [43], and the numerical stress makers are removed to reduce the phoneme set, which is shown to be beneficial for the performance [4, 28, 17]. As for Multilingual LibriSpeech, we use Phonemizer [6], [^5] https://github.com/bootphon/phonemizer a toolkit that supports a wide variety of languages other than English. The language-switching labels are removed as well as the phonemes appear less than 1000 times in the text data. For the Italian, we isolate the double consonant symbol in the lexicon, which is crucial for the initial stage to converge. Note that for TIMIT we do not adopt any of the above data preparations since it originally contains its phone inventories including silence. We use the 39-phone inventory which can be easily found in Kaldi [46] for training and evaluation, [^6] https://github.com/kaldi-asr/kaldi/blob/master/egs/timit/s5/conf/phones.60-48-39.map following prior UASR works [31, 4, 17].
@@
After the data preparation, we directly use the phonemicized text with <silence> token to train the phoneme LM. More specifically, we utilize KenLM [21], [^7] https://github.com/kpu/kenlm to build the 4-gram LM mentioned in 3.1.3. Since in the GAN training stage, we are actually matching the two distributions between the segmented speech features and phonemicized text with <silence>, we naturally build our external LM for RL segmentation learning from the same preprocessed text. The LM is also utilized for the unsupervised cross-validation described in E.5.
E.2 Self-Training with Hidden Markov Models
@@
Originated from semi-supervised learning, self-training aims at providing a trivial approach to utilize the unpaired or unlabeled data effectively [56, 33, 23]. More specifically, given a supervised ASR system $S_{ASR}$ learned from the labeled speech-text data $D_{l}$, we can use $S_{ASR}$ to generate pseudo labels of a large but unlabeled speech data $D_{s}$. Now, we can utilize the speech data along with the pseudo labels to perform supervised learning or re-tuning [56] to obtain a new ASR system $S^{\prime}_{ASR}$. Generally, given that the pseudo label is good enough, we may expect that $S^{\prime}_{ASR}$ performs better than $S_{ASR}$.
@@
In the unsupervised setup, the core concept of self-training is suited as well. By simply using an unsupervised ASR system $S_{UASR}$ to perform pseudo labeling, we can train a new ASR model $S^{\prime}_{UASR}$ by the supervised objective. Since no labeled data is used (we are just using the pseudo labels), the new ASR system does not violate the unsupervised setup. More specifically, we use Hidden Markov Models (HMMs; Bosch et al. [7])as the backbone of the new ASR system. This can be dated back to [10] and is also adopted by recent UASR [4, 28]. Furthermore, to give fair comparisons and increase reproducibility, we follow the publicly available code provided by wav2vec-U for HMM self-training. [^8] https://github.com/facebookresearch/fairseq/tree/main/examples/wav2vec/unsupervised/kaldi_self_train Our results in Table 1 indicate that with the high-quality pseudo labels generated by Reborn, the new ASR model based on HMMs performs better than the original UASR system and is the best among all the other methods.
E.3 Segmentation Model Architecture
@@
We parametrize the segmentation model $\pi_{\theta}$ using a two-layered 1-dimensional convolutional neural network (CNN). This choice of architecture allows us to efficiently extract relevant information from adjacent speech features, making it particularly well-suited for our segmentation task. The three convolutional layers in our CNN have different kernel sizes, namely 7 and 3. This is a lightweight model, making it computationally efficient for applications.
E.4 Hyperparameter Search
@@
We search the optimal hyperparameters of our phoneme predictor and segmentation models with unsupervised validation metrics. For the phoneme predictor, we directly adopt the search space of the hyperparameters and the unsupervised validation metric from wav2vec-U [4].We search the hyperparameters indicated as $\lambda$, $\gamma$, and $\eta$ in Appendix D only during the initialization stage of the phoneme predictor (§ 3.3). We directly adopt the same hyperparameters in the following iterations of GAN training. As for the segmentation model, we take the idea from Baevski et al. [4] and derive a similar unsupervised metric.
E.5 Unsupervised Validation Metric for Segmentation Model Selection
@@
$\theta^{*}=\arg\min_{\theta}\left(\frac{{-\sum_{i=1}^{M}\text{ log }p_{LM}(y^{\prime}_{i,\theta})}}{{U(Y^{\prime}_{\theta})}}\right)$ (10)
E.6 Training Details
@@
E.7 WFST Decoding for Obtaining Word-Level Outputs
@@
Appendix F Speech Self-supervised Learning and Acoustic Unit Discovery
Self-supervised learning (SSL) aims to explore effective ways for representation learning without labeled data. Recent speech SSL models can be categorized into the three classes based on the pre-training objective [35], including generative [29, 27, 30], contrastive [40, 2, 3], and predictive [22, 5, 12]. The SSL representation can be utilized for many different downstream tasks, such as phoneme recognition, speaker identification, or emotion recognition [58]. Recent UASR also takes advantage from speech representations from SSL models [4, 17].
@@ Appendix F Speech Self-supervised Learning and Acoustic Unit Discovery
Appendix G Different Schemes for Boundary Evaluation

### v3

@@ A.2 Reward Design Details
A.3 Iteratively Polished Phoneme Predictors
@@
A.4 Generalizability across Different Speech Foundation Models
@@
$\mathcal{L}_{\text{GAN}}=$ $\min_{\mathcal{G}}\max_{\mathcal{D}}\underset{Z\sim\mathcal{Z}}{\mathbb{E}}[\log(\mathcal{D}(Z))]+\underset{S\sim\mathcal{S}}{\mathbb{E}}[\log(1-\mathcal{D}(\mathcal{G}(S)))]$ (5)
, where $Z$ is a tokenized phoneme sequence sampled from the phonemicized text corpora $\mathcal{Z}$ and $S$ is the preprocessed speech feature sampled from the preprocessed speech corpora $\mathcal{S}$. Besides the min-max objective in the adversarial training, wav2vec-U [4] also adopts the three regularized objectives for better convergence and performance. First, the gradient penalty $\mathcal{L}_{\text{gp}}$ [21] is applied on the discriminator $\mathcal{D}$ to stabilize training, where the input of $\mathcal{D}$ is the linear combination of a generated sample $G$ and a real sample $Z$.
$\mathcal{L}_{\text{gp}}=\underset{G,Z,\alpha\sim U(0,1)}{\mathbb{E}}\left[\left(\left\|\nabla\mathcal{D}(\alpha G+(1-\alpha)Z)\right\|_{2}-1\right)^{2}\right]$ (6)
Next, the smoothness penalty $\mathcal{L}_{\text{sp}}$ is computed across consecutive segments of a generated logit sequence, promoting the generation of closely aligned neighboring outputs by the model.
$\mathcal{L}_{\text{sp}}=\sum_{(g_{t},g_{t+1})\in G}\left\|g_{t+1}-g_{t}\right\|^{2}$ (7)
Furthermore, the phoneme diversity loss $\mathcal{L}_{\text{pd}}$ is applied to encourage higher vocabulary usage of the generated sequences. As shown in Eq. 8, the entropy of the averaged softmax distribution of the generator outputs ($H_{\mathcal{G}}(\mathcal{G}(S))$) is maximized over the phoneme vocabulary across a batch B.
$\mathcal{L}_{\text{pd}}=\frac{1}{|B|}\sum_{S\in B}-H_{\mathcal{G}}(\mathcal{G}(S))$ (8)
@@
$\mathcal{L}=\mathcal{L}_{\text{GAN}}-\lambda\mathcal{L}_{\text{gp}}+\gamma\mathcal{L}_{\text{sp}}+\eta\mathcal{L}_{\text{pd}}$ (9)
@@
Appendix C Implementation Details
C.1 Data Preparation and Phoneme LM Formulation
@@
Our data preparation process mainly follows wav2vec-U [^1] https://github.com/facebookresearch/fairseq/tree/main/examples/wav2vec/unsupervised for fair comparisons. Specifically, we use wav2vec 2.0 large pretrained on LibriLight [24] as the feature extractor in English [^2] https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_new.pt; and the multilingual version of wav2vec 2.0 (XLSR-53) for the other langauges [^3] https://dl.fbaipublicfiles.com/fairseq/wav2vec/xlsr_53_56k.pt. Before the feature extraction stage, we attempt to remove most of the silence through an unsupervised voice activity detection (VAD) method [58] [^4] https://github.com/zhenghuatan/rVAD. However, the method is not completely perfect. To stabilize the GAN training, a special <silence> token is introduced to the text data. The <silence> token is inserted in the front and the end of each sentence, and also between words by a probability of 0.25 according to wav2vec-U. Then, we phonemicize all the words in the text data along with the previously inserted <silence> tokens for GAN training. The two off-the-shelf phonemizers are used to perform phonemicization. For LibriSpeech, we use the G2P phonemeizer [46], and the numerical stress makers are removed to reduce the phoneme set, which is shown to be beneficial for the performance [4, 30, 18]. As for Multilingual LibriSpeech, we use Phonemizer [6] [^5] https://github.com/bootphon/phonemizer, a toolkit that supports a wide variety of languages other than English. The language-switching labels are removed as well as the phonemes appear less than 1000 times in the text data. For the Italian, we isolate the double consonant symbol in the lexicon, which is crucial for the initial stage to converge. Note that for TIMIT we do not adopt any of the above data preparations since it originally contains its phone inventories including silence. We use the 39-phone inventory which can be easily found in Kaldi [49] for training and evaluation [^6] https://github.com/kaldi-asr/kaldi/blob/master/egs/timit/s5/conf/phones.60-48-39.map, following prior UASR works [33, 4, 18].
@@
After the data preparation, we directly use the phonemicized text with <silence> token to train the phoneme LM. More specifically, we utilize KenLM [22] [^7] https://github.com/kpu/kenlm, to build the 4-gram LM mentioned in 3.1.3. Since in the GAN training stage, we are actually matching the two distributions between the segmented speech features and phonemicized text with <silence>, we naturally build our external LM for RL segmentation learning from the same preprocessed text. The LM is also utilized for the unsupervised cross-validation described in C.5.
C.2 Self-Training with Hidden Markov Models
@@
Originated from semi-supervised learning, self-training aims at providing a trivial approach to utilize the unpaired or unlabeled data effectively [59, 35, 25]. More specifically, given a supervised ASR system $\mathrm{M}_{\text{ASR}}$ learned from the labeled speech-text data $\mathrm{D}_{\text{labeled}}$, we can use $\mathrm{M}_{\text{ASR}}$ to generate pseudo labels of a significant but unlabeled speech data $\mathrm{D}_{\text{unlabeled}}$. Now, we can utilize the speech data along with the pseudo labels to perform supervised learning or re-tuning [59] to obtain a new ASR system $\mathrm{M^{\prime}}_{\text{ASR}}$. Generally, given that the pseudo label is good enough, we may expect that $\mathrm{M^{\prime}}_{\text{ASR}}$ performs better than $\mathrm{M}_{\text{ASR}}$.
@@
In the unsupervised setup, the core concept of self-training is suited as well. By simply using an unsupervised ASR system $\mathrm{M}_{\text{UASR}}$ to perform pseudo labeling, we can train a new ASR model $\mathrm{M^{\prime}}_{\text{UASR}}$ by the supervised objective. Since no labeled data is used (we are just using the pseudo labels), the new ASR system does not violate the unsupervised setup. More specifically, we use Hidden Markov Models (HMMs; Bosch et al. [8])as the backbone of the new ASR system. This can be dated back to [11] and is also adopted by recent UASR [4, 30]. Furthermore, to give fair comparisons and increase reproducibility, we follow the publicly available code provided by wav2vec-U for HMM self-training [^8] https://github.com/facebookresearch/fairseq/tree/main/examples/wav2vec/unsupervised/kaldi_self_train. Our results in Table 1 indicate that with the high-quality pseudo labels generated by Reborn, the new ASR model based on HMMs performs better than the original UASR system and is the best among all the other methods.
C.3 Model Architecture
@@
We parametrize the segmentation model $\pi_{\theta}$ using a two-layer 1-dimensional convolutional neural network (CNN). This choice of architecture allows us to efficiently extract relevant information from adjacent speech features, making it particularly well-suited for our segmentation task. The two convolutional layers in our CNN have different kernel sizes, namely 7 and 3. This is a lightweight model, making it computationally efficient for applications. The phoneme prediction model (generator) and the discriminator in GAN training are parametrized by a single-layer 1-dimensional CNN and a two-layer 1-dimensional CNN, respectively, following wav2vec-U. Note that the phoneme prediction model generates the phoneme sequences non-autoregressively due to its architecture.
C.4 Hyperparameter Search
@@
We search the optimal hyperparameters of our phoneme predictor and segmentation models with unsupervised validation metrics. For the phoneme predictor, we directly adopt the search space of the hyperparameters and the unsupervised validation metric from wav2vec-U [4].We search the hyperparameters indicated as $\lambda$, $\gamma$, and $\eta$ in Appendix B only during the initialization stage of the phoneme predictor (§ 3.3). We directly adopt the same hyperparameters in the following iterations of GAN training. As for the segmentation model, we take the idea from Baevski et al. [4] and derive a similar unsupervised metric.
C.5 Unsupervised Validation Metric for Segmentation Model Selection
@@
$\theta^{*}=\arg\min_{\theta}\left(\frac{{-\sum_{i=1}^{M}\text{ log }p_{\text{LM}}(y^{\prime}_{i,\theta})}}{{U(Y^{\prime}_{\theta})}}\right)$ (10)
C.6 Training Details
@@
C.7 WFST Decoding for Obtaining Word-Level Outputs
@@
Appendix D Speech Self-supervised Learning and Acoustic Unit Discovery
Self-supervised learning (SSL) aims to explore effective ways for representation learning without labeled data. Recent speech SSL models can be categorized into the three classes based on the pretraining objective [37], including generative [31, 29, 32], contrastive [43, 2, 3], and predictive [23, 5, 13]. The SSL representation can be utilized for many different downstream tasks, such as phoneme recognition, speaker identification, or emotion recognition [62]. Recent UASR also takes advantage from speech representations from SSL models [4, 18].
@@ Appendix D Speech Self-supervised Learning and Acoustic Unit Discovery
Appendix E Different Schemes for Boundary Evaluation